How does F^{mu nu} tensor is connected with the E and B vectors?
1
Expert's answer
2019-06-18T08:09:00-0400

By definition 


"F^{\mu\nu}=\partial^\mu A^\nu-\partial^\nu A^\mu"

where (c=1)


"\partial^\mu=(\frac{\partial}{\partial t},-\nabla)\,,\quad
\nabla=(\frac{\partial}{\partial x},\frac{\partial}{\partial y},\frac{\partial}{\partial z})"

Connection between the vectors "\vec{E}" , "\vec{B}" and 4-vector "A^\mu=(A^0,\vec{A})" is


"\vec{E}=-\nabla A^0-\partial^0\vec{A}\,,\quad
\vec{B}=\operatorname{curl}\vec{A}=\nabla\times\vec{A}"

So

"F^{0i}=\partial^0 A^i-\partial^i A^0=\partial^0 A^i+\nabla_i A^0=-\vec{E}_i\,,\quad i=1,2,3"

"F^{12}=\partial^1 A^2-\partial^2A^1=-\partial_x A_y+\partial_y A_x=-[\nabla\times\vec{A}]_z=-{B}_z"

"F^{13}=\partial^1 A^3-\partial^3A^1=-\partial_x A_z+\partial_z A_x=[\nabla\times\vec{A}]_y={B}_y""F^{23}=\partial^2 A^3-\partial^3A^2=-\partial_y A_z+\partial_z A_y=-[\nabla\times\vec{A}]_x=-{B}_x"

Taking into account that "F^{\mu\nu}=-F^{\nu\mu}" and "F^{\mu\mu}=0" we have

"F^{\mu\nu}=
\left(
 \begin{array}{cccc}
 0&-E_x&-E_y&-E_y\\
 E_x&0&-B_z&B_y\\
 E_y&B_z&0&-B_x\\
E_z&-B_y&B_x&0\\
 \end{array}
\right)"